# Instructions for *ACL Proceedings

### v1

@@ 4.2 Main Experiment
As demonstrated in Table 2, Think-J achieved the best performance across all benchmarks, surpassing both close-sourced and fine-tuned judges. Notably, our method required only 9832 training samples, but still achieve marginal improvement on 32B-sized models. This underscores the effectiveness of Think-J, which leverages thinking optimization with the correctness of judgment as feedback to enhance preference modeling capability.
Furthermore, as Table 3 illustrates, starting from the same base model and training data, our proposed method consistently outperforms other approaches for fine-tuning LLM judges. This demonstrates the effectiveness of our judgment thinking optimization. In contrast, both naive and fine-tuned generative methods yield inferior results. Notably, the rejection sampling method, despite being trained on the same data constructed by the critic, also underperforms. This underscores the critical role of learning from negative samples when modeling human preference Liu et al. (2024b).
5 Analysis
5.1 Less is More for Thinking Initialization
We compared the performance of different data source for judgment thinking initialization in Figure 3. Our findings indicate that a small amount of thinking trace annotated with Deepseek-R1 can significantly enhance the model’s judgment capabilities. In contrast, using thinking trace annotated with Deepseek-V3, or removing the trace from the data would results in a substantial decline in performance. This highlights the importance of high-quality trace for thinking initialization.
We also compare the impact of different data selection strategies in Table 5. The results show that data quality is crucial for effective model initialization. For instance, a model initialized with chatbot-arena141414huggingface.co/datasets/lmsys/chatbot_arena_conversations, which contains relatively noisy data, achieves minimal improvement. Conversely, selecting the longest traces proves detrimental, as the longest traces are often code or math-related, which can negatively impact the data diversity.
@@ 5.1 Less is More for Thinking Initialization
Table 4: Experiment results of different RL methods.
@@
w/o clip offline 82.6 88.1 70.2 81.4
online 79.6 85.1 67.5 78.7
w/ clip offline 82.8 87.4 77.2 83.2
@@
Table 6: Experiment results of different initialization strategies for RL optimization on HelpSteer2.
We further investigated the impact of different initialization strategies on the subsequent optimization process, as shown in Table 6. The results indicate that even without thinking initialization, RL-based methods can still achieve substantial improvements, validating their effectiveness. Moreover, initializing the model with a few R1-annotated traces leads to a more structured and effective reasoning pattern, resulting in further enhancements in performance. In contrast, when initialization is performed without trace clipping, the optimization effect deteriorates, particularly for online RL. This finding supports our earlier hypothesis that for general LLM judgment, longer thinking do not necessarily correlate with better accuracy.
5.2 Best Practice for Thinking Optimization
@@
Setting Qwen2.5-32B-Instruct Meta-Llama-3-8B-Instruct
RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
$\beta$=1.0 96.7 83.2 90.1 92.0 90.5 93.9 74.3 90.3 77.8 84.1
$\beta$=0.5 96.9 85.8 91.2 87.4 90.3 95.0 73.3 87.4 77.2 83.2
$\beta$=0.0 92.9 71.3 86.3 67.3 79.4 96.2 65.5 88.0 71.5 80.3
$\gamma$=0.5 91.2 75.2 71.2 61.5 74.8 82.7 65.5 80.4 66.4 73.7
$\gamma$=0.2 96.7 83.2 90.1 92.0 90.5 96.5 66.8 88.7 73.2 81.3
$\gamma$=0.0 96.0 83.7 90.5 87.7 89.5 93.9 74.3 90.3 77.8 84.1
Table 8: Experiment results of different reward function settings on Helpsteer2-Preference.
With new RL algorithms continue to emerge in LLM training Zhang et al. (2025), in this section, we compare different RL methods and strategies for judgment thinking optimization.
As shown in Table 4, for offline RL, constructing traces based on sampling (with a large beam size) would lead to performance degradation. This verifies the critical role of critic model for constructing both positive and negative examples. On the other hand, for online RL methods, we find that PPO Schulman et al. (2017) significantly underperforms compared to GRPO and Reinforce++ Hu et al. (2025). This discrepancy suggests that incorporating a value model for thinking optimization would decrease training stability. We argue that natural language generation (NLG) tasks is distinct from the sequential decision-making tasks in traditional RL. Therefore, the introduction of an additional value model is not only unnecessary but may also hinder training efficiency.
Finally, perform offline learning in an iterative manner, as shown in Table 7, can achieve further improvement. However, this will result in a more complex and cumbersome training pipeline.
@@ 5.3 Reward Design for Online Learning
In this section, we aimi to analyze the contribution of different components of the reward function as defined in 6. We fixed $\alpha$ as 1.0 and varied the weights $\beta$, and $\gamma$, with the results detailed in Table 8. Our findings indicate that the formatting reward $r_{\text{format}}$ is crucial for both models, as its removal leads to significant performance degradation. Conversely, the influence of the strength reward $r_{\text{strength}}$ differs between the models. For Llama3-8B, removing $r_{\text{strength}}$ results in a performance improvement, while it offers a slight enhancement for the more capable Qwen2.5-32B. We hypothesize that Qwen2.5-32B’s stronger inherent abilities allow it to effectively learn the correlation between preference strength and judgment. In contrast, for the weaker model Llama-3-8B, this additional learning objective may introduce confusion.
@@ 5.4 Thinking Makes the Judge More Robust
In real applications, it is common that there exists noise in the training set, or there is a distributional difference between the training and test sets. In such cases, Think-J demonstrates superior robustness as a result of its judgment thinking ability.
To verify this, we adopted the approach from Wang et al. (2024b) and divided HH-RLHF into four groups with different data quality [^15] Data quality is indicated by the difference in scores assigned to response pairs by the reward model. For more details, please refer to Wang et al. (2024b).. We then trained judges on the data based on classifier-judge or Think-J. As shown in Figure 4, for the two groups with higher data quality, classifier-based judge achieve comparable or even better performance. However, for the groups with lower quality, the accuracy of classifier-based methods drops significantly, even falling below random guessing. In contrast, Think-J maintains relative stability, verifying its robustness to varied data quality.
@@ 6 Conclusion
In this paper, we propose Think-J to enhance generative LLM-as-Judge with judgment thinking optimization. Experiment results verify the effectiveness of Think-J compared with both classifier-based and generative LLM judges.
With the increasing popularity of RL-based test-time scaling methods, it is crucial to develop a reliable and stable feedback system that aligns well with real-world human preferences. In future, we will continue to explore generative judges for more accurate preference modeling and optimization.
Limitations
Our work still has several limitations: 1) Due to time and resource constraints, we only validated our method on 32B-sized models with a relatively small dataset. Fine-tuning larger models with more comprehensive preference data warrants future investigation. 2) While LLM-as-a-Judge and Reward Modeling are two inter-correlated tasks, the lack of absolute scoring data currently prevents us from extending our pairwise-trained LLM-Judge to a pointwise setting, hindering it for direct RL guidance. Acquiring such data is a key direction for future work. 3) The increasing prevalence of thinking-enhanced models like Qwen-3 necessitates investigating the continued effectiveness of our method on these architectures. Scaling our approach to reasoning-enhanced models would further strengthen our findings.
Ethical Considerations
Our research aims to model human preference, ultimately making LLM outputs less harmful and more helpful. While the judge model might occasionally produce relatively harmful outputs during experimentation, its core purpose is to provide unbiased judgment towards helpfulness and harmlessness, thereby reducing the occurrence of such outputs. Furthermore, our method leverages existing datasets rather than creating new ones. Based on these points, we believe our work is ethical.
@@ A.1 Margin-based Reward for GRPO
As we have explained in Section 3.2.2, we design the reward function as the combination of three parts: $r_{\text{accuracy}}$, $r_{\text{format}}$, $r_{\text{strength}}$. Referring to the work of Chen et al. (2025a), for the purpose of more fine-grained judgment prediction, we also tested the following reward function design:
$r_{\text{margin}}$ =$=\begin{cases}-||\text{s}_{\text{resp1}}-\text{s}_{\text{resp2}}||,&\text{if }\text{judgement}=\text{label}\\ ||\text{s}_{\text{resp1}}-\text{s}_{\text{resp2}}||,&\text{if }\text{judgement}\neq\text{label}\end{cases}$
@@
<think> {thinking trace} <think>
@@
Table 9: Experiment results of the impact of $r_{margin}$ on Helpsteer2-Preference.
@@
A.2 Statistical Indicators of Judgment Thinking Optimization
@@
In this section, we present the detailed statistical indicators during Judgment Thinking Optimization. The statistical indicators of offline learning, GRPO training, PPO Schulman et al. (2017) training and Reinforce++ Hu et al. (2025) training are presented in Figure 5, 6, 8, 7, respectively.
@@
As the results indicate, GRPO demonstrates consistent and substantial increases in predicted rewards, a strong signal of learning to generate desirable outputs, and the critic’s low and stable value function loss suggests reliable outcome evaluation. While DPO effectively aligns with preferences, and Reinforce++ and PPO show progress in reward acquisition and stability, GRPO’s robust reward improvement combined with a well-functioning critic positions it as the most robust algorithm based on these indicators.
Notably, GRPO’s response length exhibits a consistent gradual increase, similar to prior work of Shao et al. (2024), while other methods start to fluctuate after a certain point. This suggests GRPO effectively learns to generate more detailed reasoning, enhancing judgment accuracy.
@@
Figure 5: The variation of statistical metrics during offline learning trained on Helpsteer2-Preference.
@@
Figure 6: The variation of statistical metrics during GRPO training on Helpsteer2-Preference.
@@
Figure 7: The variation of statistical metrics during Reinforce++ training on Helpsteer2-Preference.
@@
Figure 8: The variation of statistical metrics during PPO training on Helpsteer2-Preference.

### v2

@@ 4.2 Main Experiment
Table 4: Experiment results of different RL strategies.
As demonstrated in Table 2, Think-J-32B achieved the best performance across all benchmarks, surpassing both close-sourced and fine-tuned judges. Notably, our method required only 9832 training samples, but still achieve marginal improvement on 32B-sized models. This underscores the effectiveness of Think-J, which leverages thinking optimization with the correctness of judgment as feedback to enhance preference modeling capability.
Furthermore, as Table 3 illustrates, starting from the same base model and training data, our proposed method consistently outperforms other approaches for fine-tuning LLM judges. This demonstrates the effectiveness of our judgment thinking optimization. In contrast, both naive and fine-tuned generative methods yield inferior results. Notably, the rejection sampling method, despite being trained on the same data constructed by the critic, also underperforms. This underscores the critical role of learning from negative samples when modeling human preference (Liu et al. 2024b).
5 Analysis
5.1 Less is More for Thinking Initialization
We compared the performance of different data source for judgment thinking initialization in Figure 3. Our findings indicate that a small amount of thinking trace annotated with Deepseek-R1 can significantly enhance the model’s judgment capabilities. In contrast, using thinking trace annotated with Deepseek-V3, or removing the trace from the data would results in a substantial decline in performance. This highlights the importance of high-quality thinking trace for judgment thinking initialization.
We also compare the impact of different data selection strategies in Table 5. The results show that data quality is crucial for effective model initialization. For instance, a model initialized with chatbot-arena(Chiang et al. 2024), which contains relatively noisy data, achieves minimal improvement. Conversely, selecting the longest traces proves detrimental, as the longest traces are often code or math-related, which can negatively impact the data diversity.
@@ 5.1 Less is More for Thinking Initialization
w/ init offline 82.8 87.4 77.2 83.2
@@
Table 6: Experiment results of different initialization strategies for RL optimization on HelpSteer2-Pref.
Setting Qwen2.5-32B-Instruct Qwen2.5-7B-Instruct
RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
$\beta$=1.0 96.7 83.2 90.1 92.0 90.5 96.1 78.6 85.9 80.4 85.3
$\beta$=0.5 96.9 85.8 91.2 87.4 90.3 95.0 77.3 87.6 77.2 84.3
$\beta$=0.0 92.9 71.3 86.3 67.3 79.4 95.2 71.5 88.0 71.7 81.6
$\gamma$=0.5 91.2 75.2 71.2 61.5 74.8 93.6 63.4 82.0 74.1 78.3
$\gamma$=0.2 96.7 83.2 90.1 92.0 90.5 94.7 76.3 88.7 78.2 84.5
$\gamma$=0.0 96.0 83.7 90.5 87.7 89.5 96.1 78.6 85.9 80.4 85.3
Table 7: Experiment results of different reward function settings on Helpsteer2-Pref.
@@
We further investigated the impact of different initialization strategies on the subsequent optimization process, as shown in Table 6. The results indicate that even without thinking initialization, RL-based methods can still achieve substantial improvements, validating their effectiveness. Moreover, initializing the model with a few R1-annotated traces leads to a more structured and effective reasoning pattern, resulting in further enhancements in performance.
5.2 Best Practice for Thinking Optimization
With new RL algorithms continue to emerge in LLM training (Zhang et al. 2025), in this section, we compare different RL strategies for judgment thinking optimization.
As shown in Table 4, for offline learning, removing SFT term will lead to performance degradation, as SFT is conducted on token-level and can provide regularization for better preference optimization. On the other hand, for online learning, we find that PPO (Schulman et al. 2017) significantly underperforms compared to GRPO and Reinforce++ (Hu et al. 2025). This discrepancy suggests that incorporating a value model for thinking optimization would decrease training stability. We argue that natural language generation (NLG) tasks is distinct from the sequential decision-making tasks in traditional RL. Therefore, the introduction of an additional value model is not only unnecessary but may also hinder training efficiency.
Finally, as shown in Table 8, offline learning in an iterative manner can achieve further improvement. However, this will result in a more complex and cumbersome training pipeline. To draw a conclusion, the best practice for thinking optimization is GRPO with carefully designed rewards.
@@ 5.3 Reward Design for Online Learning
In this section, we aim to analyze the contribution of different components of the reward function as defined in 6. We fixed $\alpha$ as 1.0 and varied the weights $\beta$, and $\gamma$.
As shown in the results in Table 7, the formatting reward $r_{\text{format}}$ is crucial for both models, as its removal leads to significant performance degradation. Conversely, the influence of the strength reward $r_{\text{strength}}$ differs between the models. For Qwen-2.5-7B-Inst, removing $r_{\text{strength}}$ results in a performance improvement, while it offers a slight enhancement for the more capable 32B model. We hypothesize that 32B model’s stronger inherent abilities allow it to effectively learn the correlation between preference strength and judgment. In contrast, for the weaker 7B model, this additional learning objective may introduce confusion.
@@ 5.4 Thinking Makes the Judge More Robust
In real applications, it is common that there exists noise in the training set, or there is a distributional difference between the training and test sets. In such cases, Think-J demonstrates superior robustness compared with classifiers as a result of its judgment thinking ability.
To verify this, we adopted the approach from (Wang et al. 2024b) and divided HH-RLHF into four groups with different data quality [^11] Data quality is indicated by the difference in scores assigned to response pairs by an external reward model. For more details please refer to the work of (Wang et al. 2024b).. We then trained judges on the data based on classifier-judge or Think-J. As shown in Figure 4, for the two groups with higher data quality, classifier-based judge achieve comparable or even better performance. However, for the groups with lower quality, the accuracy of classifier-based methods drops significantly, even falling below random guessing. In contrast, Think-J maintains relative stability, verifying its robustness to varied data quality [^12] We also present thinking trace error analysis in Appendix A.6..
@@ 6 Conclusion
In this paper, we propose Think-J to enhance generative LLM-Judges with judgment thinking optimization. Experiment results verify the effectiveness of Think-J compared with both classifier-based and generative LLM judges. With the increasing popularity of RL-based test-time scaling methods, it is crucial to develop a reliable and stable feedback system that aligns well with real-world human preferences. In future, we will continue to explore generative judges for more accurate preference modeling.
Acknowledgements
This work was supported in part by the Jiangsu Science and Technology Major Project (BG2024031) and Nanjing University AI & AI for Science Funding (2024300540).
@@ A.1 Margin-based Reward for GRPO
As we have explained in Section 3.2, we design the reward function as the combination of three parts: $r_{\text{accuracy}}$, $r_{\text{format}}$, $r_{\text{strength}}$. Referring to the work of Chen et al., for the purpose of more fine-grained judgment prediction, we also tested the following reward function design:
[…]
@@
<think> {thinking trace} </think>
@@
Table 9: Experiment results of the impact of $r_{margin}$ on Helpsteer2-Pref.
@@
A.2 Efficiency of Trace Clipping
As we explained in Section 3.1 and Appendix B.1, we perform trace clipping for the thinking traces generated by Deepseek-R1 to reduce inference overhead and improve traning efficiency. To verify this, we present the results of thinking trace initialization with and without trace clipping in Table 10, which clearly shows that when initialization is performed without trace clipping, the optimization effect deteriorates, particularly for online RL. This finding supports our earlier hypothesis that for general LLM judgment, longer thinking do not necessarily correlate with better accuracy.
Init RL RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
w/o clipping offline 82.6 88.1 70.2 81.4
online 79.6 85.1 67.5 78.7
w/ clipping offline 82.8 87.4 77.2 83.2
online 82.9 90.3 77.8 84.1
Table 10: Comparison of thinking trace initialization with or without trace clipping.
Moreover, beyond these performance gains, trace clipping also significantly reduces inference time and computational requirements. To quantify this, we evaluated two online-optimized Qwen-2.5-7B-Instruct models from Table 6 on a single A100-80G GPU. This evaluation involved bidirectional inference across 5,970 samples from RewardBench, and we measured both total inference time and average thinking trace length (tokenized by the Qwen2.5-7B-Instruct tokenizer). As evident from Table 11, trace clipping drastically reduced inference time, primarily due to the considerable decrease in thinking trace length.
In conclusion, by fostering a more concise thinking process for judgment thinking initialization, trace-clipping not only accelerates inference but also enhances optimization performance.
Init RL Inference Time(s) Avg Trace Length
w/o clip online 2947.01 907.06
w/ clip online 785.42 231.26
Table 11: Comparison of inference efficiency of trace clipping on Qwen-2.5-7B-Instruct.
A.3 Critic-model for Preference Pair Construction
While some of previous construct preference pairs for DPO by sampling multiple responses, In Section 3.2, we primarily employ a critic model for constructing the preference pairs. This is because, after judgment thinking initialization, the model already possesses strong judging capabilities. This means simply sampling multiple times often doesn’t guarantee the creation of distinct negative examples for much of the preference data.
To verify this, we compared both the efficiency and the performance of our critic-based method and the sampling-based approach on Llama-3-8B-Instruct. Specifically, for each instruction, we sampled 16 judgments, randomly selecting one correct judgment as the positive sample and one incorrect judgment as the negative sample. If a pair of both correct and incorrect judgments couldn’t be found, the sample was dropped.
Method Original Sampling-based Critic-based
Pair Num 9125 5111 8942
Table 12: Comparison of preference sample number created by different methods on Llama-3-8B-Instruct.
Method RewardBench
Chat Hard Safety Reasoning Overall
Baseline 90.4 44.7 76.5 63.4 68.6
Sampling-based DPO 92.9 75.0 86.9 70.1 81.2
Critic-guided DPO 95.0 73.3 87.4 77.2 83.2
Table 13: Comparison of offline learning from sampling-based or critic-guided preference pairs on Llama-3-8B-Instruct.
Data RewardBench
Chat Hard Safety Reasoning Overall
Hybirated 95.0 73.3 87.4 77.2 83.2
Solely-Critic 94.9 72.6 88.1 77.6 83.3
Table 14: Comparison of training of on data constructed solely from the critic model or in a hybrid way on Llama-3-8B-Instruct.
As shown in Table 12, for the Helpsteer2-Pref dataset, even with n=16 samples, the sampling-based method failed to generate negative samples for nearly half of the data. This significantly reduces data utilization efficiency. Therefore, our critic-based method substantially outperforms the sampling-based method, as demonstrated in Table 13.
Moreover, as we opt to using the critic model to only generate the contrast samples for the preference pairs, we also verified the effectiveness of training solely on data generated by the critic. As shown in Table 14, there was no significant difference in performance. Therefore, we decided to use the critic to construct either positive or negative samples for a given response. This approach effectively reduces inference time by half compared to constructing both.
A.4 Statistical Indicators of Judgment Thinking Optimization
@@
In this section, we present the detailed statistical indicators during Judgment Thinking Optimization. The statistical indicators of offline learning, GRPO training, PPO (Schulman et al. 2017) training and Reinforce++ (Hu et al. 2025) training are presented in Figure 5, 6, 8, 7, respectively.
@@
As the results indicate, GRPO demonstrates consistent and substantial increases in predicted rewards, a strong signal of learning to generate desirable outputs, and the critic’s low and stable value function loss suggests reliable outcome evaluation. While DPO effectively aligns with preferences, and Reinforce++ and PPO show progress in reward acquisition and stability, GRPO’s robust reward improvement combined with a well-functioning critic positions it as the most robust algorithm based on these indicators. Notably, GRPO’s response length exhibits a consistent gradual increase, similar to prior work of (Shao et al. 2024), while other methods start to fluctuate after a certain point. This suggests GRPO effectively learns to generate more detailed reasoning, enhancing judgment accuracy.
@@
Figure 5: The variation of statistical metrics during offline learning trained on Helpsteer2-Pref.
@@
Figure 6: The variation of statistical metrics during GRPO training on Helpsteer2-Pref.
@@
Figure 7: The variation of statistical metrics during Reinforce++ training on Helpsteer2-Pref.
@@
Figure 8: The variation of statistical metrics during PPO training on Helpsteer2-Pref.
A.5 Human Evaluation of Thinking Trace Interpretability
To quantitatively verify the improvement of thinking trace interpretability introduced by our method, we conducted a human evaluation. We randomly sampled 100 cases from RewardBench and enlisted three graduate students to manually assess the thinking trace produced by three judge models: the original model, the model fine-tuned with Supervised Fine-Tuning (SFT) on positive data only, and our model optimized with online Reinforcement Learning (RL), Think-J. The students were instructed to compare the two judgments given by different models for the same sample and determine which one provided a better thinking. We then recorded the majority vote from the three students.
Setting Win Rate
Think-J v.s. Original Qwen-2.5-7B-Inst 100%
Think-J v.s. SFT on Positive Samples 67%
Table 15: Comparison of interpretability of thinking traces generated by different judges.
As the results show in Table 15, Think-J unquestionably outperforms the judgments generated by the original model. This is expected, as the original model was not optimized for thinking-enhanced judgment, and its interpretations were too simplistic. Furthermore, Think-J also surpasses the judgments from the model SFT on positive samples only, underscoring the necessity of negative samples when optimizing judgment thinking traces.
A.6 Limited Failure Analysis of Think-J Robustness
While we generally assume a correct thinking trace would lead to a correct judgment (and vice versa), we acknowledge cases where an incorrect thinking trace might still produce the right outcome. To perform a more in-depth evaluation of the model’s underlying thinking process, we leveraged an LLM-as-a-Judge approach based on the data and model in Section 5.4.
Model Error Rate
Chat Hard Safety Reasoning
Think-J-Qwen-2.5 trained on Group1 32.5 55.0 60.0 63.1
Think-J-Qwen-2.5 trained on Group4 23.1 33.3 46.7 35.0
Table 16: Error rate of thinking traces generated by different Think-J models trained on different qualities groups of data.
Specifically, we randomly selected 50 samples from RewardBench and employed GPT-4o to meticulously check for errors in the judgment process (i.e., the thinking trace) for each model. We removed the final binary judgment (i.e., ”Response A/B is better”) and focused solely on verifying errors within the clipped critique itself, even when the final judgment result was correct. This methodology allowed us to better assess the robustness of Think-J’s internal logic.
Unexpectedly, as shown in Table 16, our analysis revealed a significant finding: Despite the high accuracy of the model trained on noised Group 1 data, its thinking trace quality appeared to degrade substantially, with over 50% of the traces containing identifiable errors. We hypothesize that for a binary classification task like judgment, an imperfect thinking trace can still lead to a correct judgment in many cases, provided the critical point with decisive influence is identified.